Operations Report – February 17th

Crew 222 Operations Report 02/17/2020

SOL: 1

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: N/A

Notes on non-nominal systems: N/A

Generator: yes
Hours run: 11 hours
From what time last night: 7pm
To what time this morning: 6am
List any additional daytime hours when the generator was run: N/A

Solar—SOC 93% (Before generator is run at night)

Diesel Reading – 78%

Station Propane Reading – 41%

Ethanol Free Gasoline – 0.5 gallons

Water (loft tank): 46 gallons

Water Meter: 149275.24 units

Water (static tank) – 441 gallons

Static to Loft Pump used – yes

Water in GreenHab – 138 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: yes
Hours: 143.7
Beginning charge: 100
Ending charge: 100
Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: N/A
Beginning charge: N/A
Ending charge: N/A
Currently charging: N/A

Spirit rover used: yes
Hours: 131.7
Beginning charge: 100
Ending charge:96
Currently charging: yes

Opportunity rover used: no
Hours: 69.9
Beginning charge: N/A
Ending charge: N/A
Currently charging: yes

Curiosity rover used: yes
Hours: 137.9
Beginning charge: 100
Ending charge: 91
Currently charging: yes

Notes on rovers: N/A

ATV’s Used: N/A
Reason for use: N/A
Oil Added? no
ATV Fuel Used: N/A
Hours the ATVs were Used today: 0
Notes on ATVs: N/A

HabCar used and why, where? N/A

CrewCar used and why, where? N/A

General notes and comments: N/A

Summary of internet: used for the com

Summary of suits and radios: Suits 7, 8, 9, 10 used and radio tested for the first EVA. All ok.

Summary of Hab operations: N/A

Summary of GreenHab operations: N/A

Summary of ScienceDome operations: N/A

Summary of RAM operations: Repaired the hygrometric captor of the weather station (finished). Working on the experiment of the backpack (not fully done).

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns and requests to Mission Support: N/A